The random variable Xhas a probability density function,.f{x) where

f(x)= "\begin{Bmatrix}
 a-5+x; 5<x<=6 \\
 0 ; otherwise 
\end{Bmatrix}" {

(a) Show that a=1/2

(b) Calculate P( X > 5. 7) .

(c) Find E (X), Var (X) and Var(3+X).


1
Expert's answer
2021-07-12T18:17:00-0400

Solution:

f(x)= "\begin{Bmatrix}
 a-5+x; 5<x<=6 \\
 0 ; otherwise 
\end{Bmatrix}"

(a) Since, f(x) is pdf, "\int_5^6f(x)dx=1"

"\int_5^6[a-5+x]dx=1
\\ \Rightarrow [ax-5x+\dfrac{x^2}2]_5^6=1
\\ \Rightarrow [6a-30+18]-[5a-25+12.5]=1
\\ \Rightarrow a-5+5.5=1
\\ \Rightarrow a=0.5"

(b):

Now, f(x)= "\begin{Bmatrix}
 x-4.5; 5<x<=6 \\
 0 ; otherwise 
\end{Bmatrix}"

"P(X>5.7)=\int_{5.7}^6(x-4.5)dx
\\=(\dfrac{x^2}2-4x)_{5.7}^6
\\=(18-24)-(16.245-22.8)
\\=0.555"

(c):

"E[X]=\int_5^6x(x-4.5)dx
\\=\int_5^6(x^2-4.5x)dx
\\=(\dfrac{x^3}3-4.5\times\dfrac{x^2}2)_5^6
\\=(\dfrac{x^3}3-2.25x^2)_5^6
\\=(\dfrac{216}3-2.25(36))-(\dfrac{125}3-2.25(25))
\\=5\dfrac{7}{12}"

"E[X^2]=\int_5^6x^2(x-4.5)dx
\\=\int_5^6(x^3-4.5x^2)dx
\\=(\dfrac{x^4}4-4.5\times\dfrac{x^3}3)_5^6
\\=(\dfrac{x^4}4-1.5x^3)_5^6
\\=(\dfrac{1296}4-1.5(216))-(\dfrac{625}4-1.5(125))
\\=31.25"

Now, "Var[X]=E[X^2]-(E[X])^2=31.25-(5\dfrac7{12})^2=\dfrac{11}{144}"

And "Var[3+X]=0+Var[X]=\dfrac{11}{144}"